OpenStudy (anonymous):

for what value(s) of k will the function h(x) = 4x^2 -3x +2kx +1 have two zeros? I got +/- 4 but i think its wrong so just need some clarification please.

OpenStudy (bahrom7893):

THIS IS EASSYYYY!!!!

OpenStudy (bahrom7893):

h(x) has 2 zeroes when the discriminant is positive

OpenStudy (anonymous):

0.o

OpenStudy (bahrom7893):

h(x) = 4x^2 + (-3+2k)x + 1

OpenStudy (anonymous):

I'm not stupid i know that too

OpenStudy (bahrom7893):

D = b^2 - 4*a*c D = (-3+2k) - 4*4*1 > 0 -3 + 2k - 16 > 0 -19 + 2k > 0

OpenStudy (bahrom7893):

2k > 19 k > 19/2

OpenStudy (anonymous):

Noo... its b^2 so (-3 + 2k) ^2 right?

OpenStudy (bahrom7893):

awww shoot

OpenStudy (anonymous):

:)

OpenStudy (bahrom7893):

Thanks for catching the mistake

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

So then was i right?

OpenStudy (bahrom7893):

D = (-3+2k)^2 - 4*4*1 = 9 -12k + 4k^2 - 16 = 4k^2 -12k - 7 >0

OpenStudy (anonymous):

But then what does k have to be?

OpenStudy (bahrom7893):

(2k-7)(2k+1)>0

OpenStudy (anonymous):

wait but where'd you get 9-12k

OpenStudy (bahrom7893):

(-3+2k)^2 - 4*4*1 = (9 -12k + 4k^2) - 16 = (2k-7)(2k+1)>0

OpenStudy (anonymous):

Ohhh, lol thank you!

OpenStudy (anonymous):

so then k has to be either 7/2 or -1/2 right?

OpenStudy (bahrom7893):

well when is a product of two numbers positive?

OpenStudy (anonymous):

when the numbers are negative?

OpenStudy (bahrom7893):

or?

OpenStudy (anonymous):

don't know

OpenStudy (bahrom7893):

when both numbers have the same sign: 7*1 = 7 (-7)(-1) = 7

OpenStudy (anonymous):

oh OR they're positive?

OpenStudy (bahrom7893):

OMG YOU GOT IT!!!!

OpenStudy (anonymous):

lol yay :D but that is right though right?

OpenStudy (anonymous):

Like how k has to be those two numbers?

OpenStudy (bahrom7893):

Not yet.. so it means: Case 1: (2k-7) < 0, 2k < 7, k<7/2 (2k+1) < 0, 2k<(-1), k<(-1/2) Case 2: (2k-7) > 0, 2k > 7, k > 7/2 (2k+1) > 0, 2k > -1, k > (-1/2)

OpenStudy (bahrom7893):